# RadioLLM: Introducing Large Language Model into Cognitive Radio via Hybrid Prompt and Token Reprogrammings

### v1

@@ Abstract
The increasing scarcity of spectrum resources and the rapid growth of wireless device have made efficient management of radio networks a critical challenge. Cognitive Radio Technology (CRT), when integrated with deep learning (DL), offers promising solutions for tasks such as radio signal classification (RSC), signal denoising, and spectrum allocation. However, existing DL-based CRT frameworks are often task-specific and lack scalability to diverse real-world scenarios. Meanwhile, Large Language Models (LLMs) have demonstrated exceptional generalization capabilities across multiple domains, making them a potential candidate for advancing CRT technologies. In this paper, we introduce RadioLLM, a novel framework that incorporates Hybrid Prompt and Token Reprogramming (HPTR) and a Frequency Attuned Fusion (FAF) module to enhance LLMs for CRT tasks. HPTR enables the integration of radio signal features with expert knowledge, while FAF improves the modeling of high-frequency features critical for precise signal processing. These innovations allow RadioLLM to handle diverse CRT tasks, bridging the gap between LLMs and traditional signal processing methods. Extensive empirical studies on multiple benchmark datasets demonstrate that the proposed RadioLLM achieves superior performance over current baselines.
1 Introduction
With the proliferation of wireless devices and the scarcity of spectrum resources, managing and optimizing limited wireless network resources has become increasingly challenging Nahum et al. (2023), Sathyanarayanan et al. (2023a). Integrating artificial intelligence (AI) with cognitive radio technology (CRT) offers a promising solution for optimizing spectrum sharing and enhancing communication quality Haider et al. (2015). However, traditional machine learning (ML) methods for radio cognition are limited by device noise and interference Feng et al. (2024), while deep learning (DL) approaches, driven by data, excel at tasks like radio signal classification (RSC), signal denoising, and spectrum allocation, significantly improving communication performance.
DL-based CRT frameworks simplify system complexity through end-to-end training but are often tailored to specific tasks and signal types. As shown in Figure 1, deploying these frameworks in real-world scenarios requires task-specific configurations, which limits their scalability and application in diverse industrial environments. Developing a universal CRT framework capable of handling diverse tasks and signals remains a critical challenge.
@@ 1 Introduction
The advent of LLMs has reignited interest in Artificial General Intelligence (AGI). LLMs, pretrained on large-scale datasets using autoregressive techniques, demonstrate generalization capabilities far beyond traditional models, making them promising for CRT. Despite their potential, LLMs face challenges in CRT applications, such as the need for significant computational resources and architectures tailored to radio signals.
Although LLMs have been applied to semantic communication Xie et al. (2024), network optimization Kotaru (2023), and spectrum sensing Shao et al. (2024), current methods exhibit several limitations:
1) Limited Cognitive Understanding of Radio Signals: LLMs, trained on text, lack inherent understanding of radio signals. Existing works rely on text-based prompts and external tools, limiting their ability to process raw signals. We adopt reprogramming techniques in Jin et al. (2024) to transform radio signals into formats compatible with LLM input, enabling direct signal processing and improving cognitive capabilities.
2) Inefficient Expert Knowledge Integration: Textual prompting for injecting expert knowledge often results in verbose templates with unnecessary words, increasing computational overhead Jin et al. (2024). We propose a hybrid prompting method that retrieves concise and contextually relevant tokens, reducing memory usage and runtime while preserving information richness.
3) Challenges in Capturing High-Frequency Signal Features: LLMs excel at low-frequency global information but struggle with high-frequency details critical for precise RSC tasks Bai et al. (2022), Si et al. (2022). To address this, we propose the Frequency Attuned Fusion (FAF) module, which combines high-frequency features from CNNs with the global context captured by LLMs, enhancing their performance in RSC tasks.
@@
2 RELATED WORK
2.1 Traditional DL-Based CRT Framework
Deep learning (DL)-based methods have been widely applied to critical tasks in CRT, such as RSC, signal denoising, and signal recovery. In RSC, supervised learning (SL) approaches have shown significant progress. For example, PETCGDNN used CNNs and GRUs for feature extraction to build an efficient modulation recognition model Zhang et al. (2021a). Additionally, a multi-view fusion RSC method was proposed, which leverages features from the time, frequency, and time-frequency domains to enhance performance Ke and Vikalo (2021).
The challenge of obtaining high-quality labeled data in wireless communication has driven the adoption of self-supervised learning (SSL). A Transformer-based SSL framework, TCSSAMR, was proposed for RSC Kong et al. (2023a), and MCLHN utilized masked contrastive learning with hard negatives to enhance signal diversity Xiao et al. (2024). For signal denoising and partial recovery, a Deep Denoising Network was proposed using residual learning Kaushal et al. (2016), and a time-frequency domain autoencoder was developed for denoising Chen et al. (2024).
@@ 2.1 Traditional DL-Based CRT Framework
2.2 Previous LLM-Based CRT Framework
The emergence of LLMs has brought significant advancements to AI, influencing domains such as time series analysis and CRT. For example, LLMs have been used in semantic communication systems for contextual understanding Xie et al. (2024) and in 6G edge computing for user association and resource allocation Kotaru (2023). However, these applications are mostly limited to telecommunications language understanding.
WirelessLLM has extended LLM applications to CRT tasks like power allocation and spectrum sensing, advancing AGI integration in CRT Shao et al. (2024). Yet, it relies on external tools, lacks an end-to-end signal processing pipeline, and overlooks critical tasks like WSC and signal denoising.
To overcome these challenges, we build on recent advances in time-series LLM research Jin et al. (2024), Pan et al. (2024), Cao et al. (2024) by incorporating reprogramming and hybrid prompting techniques. This enables the development of a unified and versatile CRT framework capable of handling diverse signal types and tasks.
3 Methodology
Overview: As shown in Figure 2, RadioLLM is designed with two primary components. First, the input radio signal is divided into patches to generate its signal embedding. The first component involves combining expert knowledge prompts with the top-K semantically similar anchors, retrieved from word token embeddings enriched with extensive external knowledge. These anchors and prompts are used to construct prefix prompts for the signal embedding. The second component integrates high-frequency CNN features with low-frequency signal information, feeding the fused output into the pre-trained LLM.
In this work, GPT-2 is adopted as the backbone. During training, we not only learn the mapping function between inputs and outputs but also fine-tune GPT-2 using the LoRA technique Hu et al. (2021).
@@ 3.1 Problem Statement
  • For signal reconstruction tasks (including denoising and recovery), $T=s^{\prime}(t)$, where $s^{\prime}(t)$ represents the reconstructed signal or a task-specific attribute of the original transmitted signal.
@@
3.2 Hybrid Prompt and Token Reprogramming
@@
$P^{\prime}_{t}[k,:]=E^{\prime}\left[\text{ArgTopK}\left(\max_{i=1}^{L_{T}}\gamma(P_{t}[i,:],E^{\prime}),K\right)[k],:\right],$ (3)
@@
3.3 Frequency Attuned Fusion
@@
The FAF module consists of three high-frequency extraction (HFE) layers, with the structure of each HFE layer shown in Figure. 2(C). Each HFE layer leverages convolution to detect local variations, ReLU to enhance non-linear features, and pooling to compress redundant information, effectively extracting high-frequency features from the input data. This design enables the network to better capture fine-grained details in the input.
@@
3.4 Output Projection
The features $F_{s}^{\prime}$ and $P_{t}^{\prime}$ are fed into the fine-tuned LLM module (Figure. 2 (D)) to obtain $F_{\text{llm}}$ and $P_{\text{llm}}$. After discarding the prefix $P_{\text{llm}}$, $F_{\text{llm}}$ is passed to the decoder to generate the output $O_{s}\in\mathbb{R}^{2\times L}$. In this work, we explore two decoding strategies: a linear layer for direct mapping and a shallow Transformer decoder that uses self-attention to capture complex dependencies, yielding better reconstruction. The pretraining objective is to minimize the mean squared error (MSE) loss between $O_{s}$ and the ground truth $s^{\prime}(t)$.
3.5 Training Strategy
The overall pretraining pipeline of RadioLLM and its application to downstream tasks are summarized in Algorithm 1. The process begins with the configuration of hyperparameters, including the network structure, optimizer, dataset parameters, and early stopping criteria. During pretraining, most parameters of the LLM remain frozen, with only a subset updated via LoRA Kotaru (2023). To compute the loss across batches from different datasets within the same epoch, a loss balancing factor $b_{i}$ is introduced to mitigate the impact of differing dataset scales. Upon completion of pretraining, only the non-frozen parameters are retained to reduce storage overhead.
For downstream tasks, $F^{\prime}_{s}$, extracted from the LLM output, is used as the feature representation for classification tasks. After pooling, it is fed into a linear classification head. The decoder’s output $O_{s}$ is directly used as the prediction for denoising and completing tasks.
@@ 3.5 Training Strategy
4 Experiments
4.1 Data
We used six publicly available datasets (RML2016a, RML2016b O’shea and West (2016), RML2016c O’Shea et al. (2016), RML2018a O’Shea et al. (2018), ADS-B ADS-B, and Wi-Fi Sankhe et al. (2019)) consolidated into a unified dataset. To prevent data leakage in downstream tasks, the data was split into training, validation, and test sets in a ratio of 8:1:1. For pretraining, high signal-to-noise ratio (SNR) signals ($\text{SNR}\geq 14$ dB) from the RML datasets were selected, while all signals from ADS-B and Wi-Fi datasets were included.
4.2 Implementation Details

### v3

@@ Abstract
The growing scarcity of spectrum resources and rapid proliferation of wireless devices make efficient radio network management critical. While deep learning-enhanced Cognitive Radio Technology (CRT) provides promising solutions for tasks such as radio signal classification (RSC), denoising, and spectrum allocation, existing DL-based CRT frameworks are typically task-specific and lack scalability in diverse real-world applications. This limitation naturally leads to the exploration of Large Language Models (LLMs), whose exceptional cross-domain generalization capabilities offer new potential for advancing CRT. To bridge this gap, we propose RadioLLM, a novel framework that integrates Hybrid Prompt and Token Reprogramming (HPTR) for combining radio signal features with expert knowledge, and a Frequency-Attuned Fusion (FAF) module for enhanced high-frequency feature modeling. Extensive evaluations on multiple benchmark datasets demonstrate that RadioLLM achieves superior performance compared to existing baselines in the majority of testing scenarios.
Index Terms:
Cognitive Radio Technolog, Large Language Models, Token Reprogramming
I Introduction
With the exponential proliferation of wireless devices and the escalating scarcity of spectrum resources, the efficient management and optimization of constrained wireless network resources has emerged as a pivotal challenge in modern communication systems [1, 2]. While the integration of artificial intelligence (AI) with cognitive radio technology (CRT) presents a transformative paradigm for dynamic spectrum sharing and communication quality enhancement [3], traditional machine learning (ML)-based methods for precise radio cognition are hindered by device-intrinsic noise and an increasing number of interference sources [4], rendering them effective only under specific operational conditions. In contrast, deep learning (DL) approaches demonstrate superior performance in radio signal classification (RSC), spectral denoising, and adaptive resource allocation through data-driven optimization frameworks.
Contemporary DL-based CRT architectures, though effective in reducing system complexity via end-to-end training, exhibit inherent task specificity and signal-type dependency. As illustrated in Fig 1, their deployment necessitates extensive task-specific configurations, thereby constraining operational scalability in heterogeneous industrial environments. This limitation underscores the critical need for developing universal CRT frameworks capable of handling cross-domain signal processing task.
@@ I Introduction
The emergence of LLMs has rekindled significant interest in the pursuit of Artificial General Intelligence (AGI). Pretrained on vast datasets using autoregressive learning techniques, LLMs exhibit generalization capabilities that far surpass those of conventional models, positioning them as promising candidates for application in CRT[5]. Nevertheless, their deployment in CRT contexts faces several challenges, including substantial computational demands and the need for architectures specifically adapted to the characteristics of radio signals.
Although LLMs have been employed in applications such as semantic communication [6], network optimization [7], and spectrum sensing [8, 9], existing approaches still exhibit several limitations:
1) Limited Cognitive Understanding of Radio Signals: Conventional LLMs, trained predominantly on textual data, demonstrate fundamental limitations in semantic comprehension when processing radio signals. Existing approaches in [8, 10] rely on text descriptions and external tools to transform signals into natural language prompts, which introduces two critical limitations: First, the textualization process inevitably discards essential signal features. Secondly, such methods prevent models from directly learning spatiotemporal patterns from raw signal representations. Building on [11], we employ token reprogramming to directly map raw I/Q sequences into LLM-compatible embeddings, thereby eliminating textual intermediaries and enabling true end-to-end signal cognition.
2) Inefficient Expert Knowledge Integration: Prompt engineering techniques empower LLMs with task-specific knowledge through carefully designed instructions. However, existing knowledge injection methods [11, 12] rely on fixed textual templates that incorporate expert knowledge via dataset background, task descriptions, and statistical information. This approach suffers from two critical limitations: First, excessive irrelevant lexical elements (e.g., prepositions, modifiers) in these templates consume valuable prefix token allocation. Second, the elongated prompt token sequences tend to dilute the feature representation of subsequent signal tokens.
To address these issues, we propose Hybrid Prompting, whose core innovation lies in retrieving the top-K most semantically similar anchors from software-based embedding space to replace conventional hardware-fixed text prompts. Compared with hardware-prompt approaches, our method achieves a 31.85% improvement in inference speed while maintaining a 0.85% enhancement in downstream task accuracy.
3) Challenges in Capturing High-Frequency Signal Features: The long-range attention mechanisms in LLMs excel at capturing global low-frequency signal patterns [13, 14], yet demonstrate limited sensitivity to high-frequency features such as transient pulses and phase discontinuities [15, 16]. While existing approach [9] attempts to enhance perceptual capabilities through task-specific adapter modules, these solutions fail to achieve cross-task transferability and generalization. To overcome these limitations, we propose the Frequency-Attuned Fusion (FAF) module, which synergistically integrates high-frequency local features extracted by CNNs with the global contextual representations learned by LLMs, thereby significantly improving performance in RSC tasks.
@@
II RELATED WORK
II-A Traditional DL-Based CRT Framework
DL-based methods have been widely applied to critical tasks in CRT, such as RSC, signal denoising, and signal recovery. In RSC, supervised learning (SL) approaches have shown significant progress. For example, PETCGDNN used CNNs and GRUs for feature extraction to build an efficient modulation recognition model [17]. Additionally, a multi-view fusion RSC method was proposed, which leverages features from the time, frequency, and time-frequency domains to enhance performance [18].
The challenge of obtaining high-quality labeled data in wireless communication has driven the adoption of self-supervised learning (SSL). A Transformer-based SSL framework, TCSSAMR, was proposed for RSC [19], and MCLHN utilized masked contrastive learning with hard negatives to enhance signal diversity [20]. For signal denoising and partial recovery, a Deep Denoising Network was proposed using residual learning [21], and a time-frequency domain autoencoder was developed for denoising [22].
@@ II-A Traditional DL-Based CRT Framework
II-B Previous LLM-Based CRT Framework
The emergence of LLMs has brought significant advancements to AI, influencing domains such as time series analysis and CRT. For example, LLMs have been used in semantic communication systems for contextual understanding [6] and in 6G edge computing for user association and resource allocation [7]. However, these applications are mostly limited to telecommunications language understanding.
While WirelessLLM extends LLMs to CRT tasks such as power allocation and spectrum sensing [8], it depends on external tools, lacks an end-to-end pipeline, and misses key capabilities like WSC and signal denoising. Similarly, LLM4WM, a multitask framework for channel modeling [9], employs task-specific adapters that impede the learning of unified representation and model coherence.
To address these limitations, we leverage recent advances in time-series LLMs [11], [12], [23] through model reprogramming and hybrid prompting, enabling a unified and versatile CRT framework for diverse signals and tasks.
III Methodology
Overview: As shown in Fig. 2, RadioLLM is designed with two primary components. First, the input radio signal is divided into patches to generate its signal embedding. The first component involves combining expert knowledge prompts with the top-K semantically similar anchors, retrieved from word token embeddings enriched with extensive external knowledge. These anchors and prompts are used to construct prefix prompts for the signal embedding. The second component integrates high-frequency CNN features with low-frequency signal information, feeding the fused output into the pre-trained LLM.
In this work, GPT-2 is adopted as the backbone. During training, we not only learn the mapping function between inputs and outputs but also fine-tune GPT-2 using the LoRA technique [24].
@@ III-A Problem Statement
  • For signal reconstruction tasks (including denoising and recovery), $T=\hat{s}(t)$, where $\hat{s}(t)$ represents the reconstructed signal or a task-specific attribute of the original transmitted signal.
@@
III-B Hybrid Prompt and Token Reprogramming
@@
$P^{\prime}_{t}[k,:]=E^{\prime}\left[\text{ArgTopK}\left(\max_{i=1}^{L_{T}}\gamma(P_{t}[i,:],E^{\prime}),K\right),:\right],$ (3)
@@
III-C Frequency Attuned Fusion
@@
The FAF module consists of three high-frequency extraction (HFE) layers, with the structure of each HFE layer shown in Fig. 2 (C). Each HFE layer leverages convolution to detect local variations, ReLU to enhance non-linear features, and pooling to compress redundant information, effectively extracting high-frequency features from the input data. This design enables the network to better capture fine-grained details in the input.
@@
III-D Output Projection
The features $F_{s}^{\prime}$ and $P_{t}^{\prime}$ are fed into the fine-tuned LLM module (Fig. 2 (D)) to obtain $F_{\text{llm}}$ and $P_{\text{llm}}$. After discarding the prefix $P_{\text{llm}}$, $F_{\text{llm}}$ is passed to the decoder to generate the output $O_{s}\in\mathbb{R}^{2\times L}$. In this work, we explore two decoding strategies: a linear layer for direct mapping and a shallow Transformer decoder that utilizes self-attention to capture complex dependencies, resulting in improved reconstruction. The pretraining objective is to minimize the mean squared error (MSE) loss between $O_{s}$ and the ground truth $s^{\prime}(t)$.
III-E Training Strategy
The overall pretraining pipeline of RadioLLM and its application to downstream tasks are summarized in Algorithm 1. The process begins with the configuration of hyperparameters, including the network structure, optimizer, dataset parameters, and early stopping criteria. During pretraining, most parameters of the LLM remain frozen, with only a subset updated via LoRA [7]. To ensure balanced learning when processing multiple datasets within the same epoch, a loss balancing factor $b_{i}$ is applied to normalize the gradient contributions from each dataset, which helps mitigate biases caused by differences in data scale or complexity. Upon completion of pretraining, only the non-frozen parameters are retained to reduce storage overhead.
For downstream tasks, two task-specific adaptation strategies are employed. In classification tasks, a feature representation $F^{\prime}_{s}$ is extracted from the LLM output, followed by a pooling operation, and then fed into a linear classification head. For denoising or signal reconstruction tasks, the decoder output $O_{s}$ is directly used as the final prediction.
@@ III-E Training Strategy
IV Experiments
IV-A Data Overview
In this study, seven publicly available datasets were utilized to develop and evaluate the proposed method, including RadioML2016.10a (RML16A), RadioML2016.10b (RML16B) [25], RadioML2016.04c (RML16C) [26], RadioML2022 (RML22) [27], RadioML2018.01a (RML18A) [28], the ADS-B dataset [29], and the Wi-Fi dataset [30]. To ensure fair evaluation and prevent data leakage in downstream tasks, all datasets were partitioned into training, validation, and test sets in an 8:1:1 ratio.
For pretraining, we utilized a combination of the RML16A, RML16B, RML16C, RML18A, ADS-B, and WiFi datasets. Specifically, only high-quality samples with ($\text{SNR}\geq 14\,\mathrm{dB}$) were selected from the RML series datasets to ensure robust feature learning from clean signals. For the ADS-B dataset, the entire training set was used without SNR filtering to capture the full range of real-world signal variations. For the WiFi dataset, 5% of the training samples were randomly chosen to provide representative diversity while maintaining computational efficiency. This pretraining strategy enabled the model to leverage both high-quality and diverse real-world data, thereby enhancing its generalization capability.
To fully exploit the unique characteristics of each dataset and enhance the model’s adaptability to diverse downstream tasks, we developed a modular prompt design framework. An illustrative example of our prompt template is shown in Fig. 3, where the prompt is composed of three main components: a dataset description, a task description, and input statistics.
Figure 3: An illustrative example of the prompt template used in this study. The template incorporates domain-specific descriptions, statistical attributes, and task-specific instructions, which are color-coded for clarity. This unified prompt design enables the model to process heterogeneous signal information with rich contextual and structural cues.
Specifically, for each dataset, we constructed a tailored dataset description that encapsulates its domain attributes, such as signal source, modulation types, and SNR conditions. In addition, for each pretraining objective—including denoising, recovering, and classification—we designed corresponding task-specific prompts that explicitly define the nature of the learning problem. Furthermore, in order to provide the model with fine-grained contextual cues, we incorporated input statistics into the prompt structure. These statistical descriptors (e.g., minimum, maximum, and median values, overall trend, and top lag features) are dynamically generated based on each individual data instance.
This unified and extensible prompt design enables the model to seamlessly process heterogeneous datasets and diverse learning tasks, while leveraging contextual information at both global and instance levels.
IV-B Dataset Details
RML16A: The RadioML2016.10a dataset comprises 220,000 modulated signals generated under simulated interference conditions, including carrier frequency offset and additive white Gaussian noise (AWGN). Each signal consists of 128 complex-valued I and Q samples. The dataset covers three analog modulation types—Wideband Frequency Modulation (WBFM), Amplitude Modulation Single Sideband (AM-SSB), Amplitude Modulation Double Sideband (AM-DSB)—and eight digital modulation types: Continuous Phase Frequency Shift Keying (CPFSK), Binary Phase Shift Keying (BPSK), 8-Phase Shift Keying (8PSK), Gaussian Frequency Shift Keying (GFSK), 16-Quadrature Amplitude Modulation (16QAM), 64-Quadrature Amplitude Modulation (64QAM), 4-Pulse Amplitude Modulation (4PAM), and Quadrature Phase Shift Keying (QPSK).
RML16B: The RadioML2016.10b dataset contains approximately 1.2 million modulated signals, covering the same modulation categories as RML16A except for AM-SSB. Each signal sample also consists of 128 I/Q points.
RML16C: The RadioML2016.04c dataset includes 162,060 modulated signals with the same modulation types as RML16A. Each sample similarly contains 128 I/Q data points.
RML22: The RadioML2022 dataset comprises 420,000 modulated signal samples, including eight digital modulation types—8PSK, BPSK, CPFSK, GFSK, 4PAM, 16QAM, 64QAM, and QPSK—and two analog modulation types: AM-DSB and WBFM. The signals span an SNR range from $-20\,\mathrm{dB}$ to $20\,\mathrm{dB}$, with a step size of $2\,\mathrm{dB}$.
RML18A: The RadioML2018.01a dataset serves as a large-scale benchmark for signal modulation classification, containing 2,555,904 signal samples across 24 modulation types (e.g., OOK, BPSK, QPSK, 16QAM) and 26 SNR levels from $-20\,\mathrm{dB}$ to $30\,\mathrm{dB}$ in $2\,\mathrm{dB}$ increments. Each sample consists of 1024 I/Q points, with 4096 samples per modulation type per SNR level, offering a balanced and comprehensive dataset.
ADS-B: The ADS-B dataset consists of 63,105 Automatic Dependent Surveillance–Broadcast signals captured in real-world open environments. Each signal is sampled at $50\,\mathrm{MHz}$ and contains 3000 sampling points, spanning 198 distinct signal classes.
Wi-Fi: The Wi-Fi dataset, described in [30], consists of over-the-air transmissions captured using 16 USRP X310 devices with a transmitter–receiver distance of 2 feet. The raw signals were processed into sequences of 128 complex-valued I/Q samples to maintain consistency with other datasets.
IV-C Implementation Details